OpenStudy (anonymous):

10(1.02)^(x/6)=60 what is the value of x

OpenStudy (lgbasallote):

hmmm first of all..ill ask...are you allowed to use a calculator here?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

good...first of all..divide both sides bby 10 to get rid of the ten in the left hand side...what will that give you?

OpenStudy (anonymous):

(1.02)^(x/6)=6

OpenStudy (lgbasallote):

nice...now we take the logarithm to isolate x..will you do it? :D

OpenStudy (anonymous):

I dont understand that part. What to put in the calculator

OpenStudy (lgbasallote):

no calculators yet..it's still just.. log_1.02 (6) = x/6 good?

OpenStudy (anonymous):

1.02=x?

OpenStudy (lgbasallote):

hhowd you get that?

OpenStudy (anonymous):

multiplied both sides by 1/6

OpenStudy (anonymous):

?

OpenStudy (lgbasallote):

ohh...well it doesnt work that way...one..you multiply oth sides by 6.....second..it becomes... 6log_1.02 (6) = x so...via power rule log_1.02 (6)^6 = x do you know how to input that in calculator?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

?

OpenStudy (lgbasallote):

you type log (6^6) divided by log (1.02)

OpenStudy (anonymous):

so 542.8853785?

OpenStudy (lgbasallote):

yep :D

OpenStudy (anonymous):

thank you!!!

OpenStudy (lgbasallote):

sorry for the late reps...some probs with my browser..izlagging :P haha but you're welcome <tips hat>